06-07 October 2008 On-sky TCS SW testing, M1 Collimation

Observers: D. Thompson, J.Hill (Tucson), LUCIFER folks (N.Ageorges,W.Seifert, et al)
Software Support: C.Biddick in Tucson
Telescope Operators: D.Gonzalez-Huerta, W.Wack



LUCIFER evening twilight flats?
  • Need to be open and ready to work at 5:15pm (sunset at 6:06).

Open and ready at 5:35pm. We were delayed by problems getting M3 back into the LUCIFER position and a balance issue on the telescope.

LUCIFER twilight flat data collected for filters: Br-gamma, Fe-II. Done at UT01:30

Stop and restart IIF and MCSPU at the request of SW support. MSCPU started with ./go but this gives scrolling warnings about "XMLSCHEMA_DIR". Restarted with ./gotemp, same problem. Calling Tom Sargent (UT01:42), he said to ignore this warning and proceed.

Correct initial pointing/collimation 01:52 Sending PARALLACTIC mode preset on BS9177. Failed immediately because selectAGW not run. Had to also stop and restart GCSL. Resend preset. New values: IE=-68.2 CA=-33.185

*Doug Miller called, we are discussing elongation of the stars in the LUCIFER image, even with 30s integration.

Test offsetPointing in detector XY arcsec
  • Use IRTC offset task with coord="XY"

offset 0.0 10.0 moved the star on the detector UP by ~16.8 arcsec (10 mm)
offset 0.0 -10.0 moved the source back to the original starting point
offset 0.0 -10.0 moved the source down on the detector
offset 0.0 10.0 to move back to the center
offset 10.0 0.0 moved the star on the detector LEFT (IRTC orientation)

So there is an X flip in LUCIFER with respect to the definitions for the IRTC that are currently in use.

Take set of guider data for mount filter analysis
_This test did not work because of the focus shift, see below_

Moving to BS9183, PA=250 to put the off-axis guide star in the patrol field.

  • filter=0.2 (default), take 10 minutes guiding data.
Start 19:58:01 (guiderimage1431) - 20:09:30
  • filter=0.05, take 10 minutes guiding data.
Start ... problems

Guider reported a couple of cycles of "Received invalid image" and the WFS is taking 2.25s exposures. We re-corrected pointing and re-sent the preset but the WFS images are still 2.25 seconds.

Also, now LUCIFER sees way out of focus images. We re-homed the AGW stages and re-sent the preset. It still looks bad, but the wind and humidity are rising and we are looking directly into the wind. Moving to BS9188 (23:30 +38) - EL now 70.

See wfsc000216.fits as a particularly bad example of what poor seeing can do.

  • reset filter=0.2 until analysis complete.
  • compare rms of offsets vs. filter value.

Seeing has gone really bad...the last acquisition image has ~5+ arcsec FWHM!

Checking preset reliability, as much as possible

The seeing is so bad that the WFS is having trouble and is driving the system out of collimation. We are switching

WT10_384 Y Y - 56s
WT10_149 Y N - 56s
WT10_373 Y N Y 58sT-G, 47sG-G, 44sG-A
WT10_102 Y N Y 58sT-G, 44sG-A
WT10_362 Y N Y 57sT-G, 42sG-A, 58sA-A
WT10_170 Y N Y 57sT-G, 44sG-A
WT10_170 Y Y 55sT-G (failed due to saturation)
ACT0405 Y Y 56sT-G (failed due to saturation)
WT10_089 Y N 49sT-G (peak value in adjustExposureTime is zero!)
ACT0293 Y N 59sT-G, 47sG-G

Torsten: we preset in GUIDE mode on a bright star (7th mag) right around 10:00pm MST. The acquisition image was close to saturated at ~55KDN. The returned messages were a little confusing. It reported that the exposure time was adjusted to 159ms and that this was below the threshhold of 500ms. Guiding was obviously not started (as it should not have), but we got back a "success" message.

Torsten: on another preset in GUIDE mode shortly thereafter, the preset completed but the stepfocus command sent immediately afterwards came back with an error "resumeGuiding: WFS did not resume". So it sounds like GCS did not know we did not need the WFS.

One guide mode preset failed because the acquisition image IT#1694 In transitioning from TRACK mode to GUIDE mode on the same source, presets were taking 56-58 seconds to complete. The timeline of events needs to be traced out and the overheads reduced if possible.

IT#1696 (CRITICAL!) We have had multiple instances tonight where the WFS, while running in closed loop ACTIVE mode, has reset its exposure time to 2.252s (e.g. #370 in /tmp on mu107). During this time the guide camera was integrating for only 0.5s (#3904) with a guide rate of 0.2Hz. The short WFS exposures mess up the collimation and this is a critical issue to diagnose and fix. For the AG exposure and guide rate, I would suggest scaling to the longest exposure time possible given the guide rate, without exceeding saturation limits.

Really weird problem

We has a fairly rapid transition between good collimation to having LUCIFER obviously out of focus while the WFS seemed to be running fine, although the seeing on the guide camera is poor (~2 arcsec). The problem appeared to start right after a change to 0.05 on the mount guiding filters and a stop/restart of PCS. The seeing got very bad for a while and now that it is getting better we still see this problem. We tried to remeasure a through-focus curve for LUCIFER but got inconsistent results. As a last resort we changed the mount guiding filters back to 0.2 and stop/restarted PCS again.

Really weird problem - at least partly solved. LUCIFER's internal focus stage was offset from nominal by ~263 steps. Not sure why it changed.

Note: Moving the AGw stage focus in the + direction results in a -Z4 correction

Take astrometry data on LUCIFER for orientation UT09:52

First have to correct the pointing. The guide star is on the acquisition image, but large and with a low peak brightness. Note to Torsten: perhaps an initial "guide star not found" error could be followed by a second exposure with 10x the first exposure, just to be sure.

  • Use LUCIastrom.cat:
    • Stone P @ PA=180 (RA~22:42) - not available
    • Stone A @ PA=45 (RA~00:43) - not available
    • M103 @ PA=0 (RA~01:33) - luci_20081006_0123 and _0124
    • BS9107 @ PA=270 (RA~03:32) - luci_20081006_0125 and 0126
  • Take 30s total integration per pointing

Take M1 Collimation data

  • Clear M2 active optics
  • Use collim.cat
  • Preset ACTIVE mode, PA=0, allow WFS to converge
  • Observe 20 < EL < 89 range
  • Record M1 data: X, Y, Z, RX, RY, Temp, EL
  • Make M1 collimation model, implement, test.

Target X Y Z RX RY RZ EL Temp M1Temp
?????? 0.974 1.126 -0.575 31.570 -15.110 0.00 85.2 -- 8.56
BS9107 0.749 1.304 -0.391 36.718 -12.504 0.00 81.3 10.0 -
C03+26 0.514 1.284 -0.297 38.094 -9.769 0.00 79.0 9.9 -
C03+21 1.143 0.989 0.079 40.361 -15.042 0.00 68.8 9.7 -
C03+13 1.201 0.981 0.138 41.024 -15.358 0.00 67.0 9.7 -
C03+06 1.344 0.845 0.304 40.915 -15.894 0.00 62.5 9.0 -
C03-02 1.802 0.686 0.665 40.205 -17.951 0.00 52.1 8.8 8.68
C03-09 2.029 0.763 0.945 38.760 -16.149 0.00 40.7 8.0 8.63
C03-18 2.311 0.888 1.023 37.750 -16.776 0.00 35.2 9.1 8.58
C03-28 2.220 1.710 1.209 35.408 -7.528 0.00 19.5 7.7 8.44

Note added later: JMH created a new collimation model for M1 based on this data. Both M1 and M2 were shifted together X=-1.0 mm to allow some range in X as the temperatures get colder.

Take astrometry data on LUCIFER for orientation
  • M45

Having trouble with hitting an X limit on the AGw stages. The symptom was even for an on-axis guide star the preset was returning with the error "...probe couldn't move into position". We had to home the motors to recover. Success returned as error, not warning? M1 was at its Y limit at the time (~5:30am). Cannot recover in time.

Setting up for twilight flats, AZ=270 EL=90 UT12:40.

07Oct08 DLM Moved GCS wfscimages to /Repository/AGW_Data/081007w_gcs and all GCS guiderimages to /Repository/AGW_Data/guider_081006-07

Did not get to these, maybe tomorrow?:

LUCIFER taking sky emissivity data
Can take these data in TRACK mode, seeing is not important. Preset to "blank" fields and measure count rate on sky background in the different filters.

LUCIFER morning twilight flats?

If there is time during the night:
  • Gather low-EL AGW dataset for binodal astig (need appropriate target field).
  • Re-center WFS on rotator center
  • Running stats on preset success (better done after new collim/ptg models)
    • Starting at 02:04 (after pointing model update)
    • TRACK mode (NN succeeded in MM attempts)
    • GUIDE mode (NN succeeded in MM attempts)
    • Active mode (2 succeeded in 2 attempts)
  • Usable guide star magnitude (record Rmag, FWHM, peak counts)
  • Other?

-- DavidThompson - 06 Oct 2008
Topic revision: r10 - 08 Oct 2008, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback